Friday, August 1, 2014

8 1 2014 Fidalgo Repro Check

Anacortes WA

Mid 60s to high 70s. Sunny until 2 pm then partly cloudy

Performed the anesthesia SOP with insulation on the treatment tub and ice blocks in the Treatment and recovery tubs to stabilize temperature fluctuations. Had a really poor treatment of the north sound tray and a less than great treatment of the south pop tray. Less than a third of the animals were gaping in north pop and only a little over half in the South pop were open. The Dabob tray was over 95% treated. I'm not sure we missed any brooders though as South sound only had a few and Dabob had none. I'm not sure why the animals didn't open but I assume that it might be due to the fact that they were quickly sprayed with fresh water to help knock off some of the more intense sediment deposits on them. This great to know for next time so I don't accidentally deter treatment. It seems like spawning season is winding down in Fidalgo. I'll be interested to see what the next tidal cycle does. Also yesterday I had a problem with the Hobo Shuttle. It would not read the logger pendant. I brought the logger pendant and shuttle with me to make sure neither of them was screwed up. Today I retried the pendant with the shuttle and the data offload properly. I think what happened yesterday was the shuttle go hot sitting in the sun in my tool box. The air was pretty stagnant and it got much hotter than expected. I'll double check everything on Monday to make sure the file and shuttle are in working order.

Numbers as Follow:

Temps in C

Pretreatment * Lacking the tarp and with little breeze, the trays changed temps faster than expected.
Initial  10
45      15
1.5     15

Initial     13
45         13
1.5        14
2.25      14

Initial    15
45        15
1.5       15

Salinity in ppt
Pretreatment    27
Treatment        60
Recovery         28

Brood Collection

Brood    0
Gaping   28
Dead      0
Closed   68

Brood     3
Gaping    45
Dead       0
Closed    38

#    size    sick
1    30      W
2    35      W
3    33      W

Brood   0
Gaping  84
Dead     0
Closed  2

As for Pics. There were so few treated in the North pop that the tube label and calipers are on the tile with the gapers and the rest of the tiles are closed. The South pop tray is like normal with the brooders between the tube label and calipers. The Dabob tray had so many gapers that they are on all tiles but the closed guys are above the tube and calipers.

My face when treatments don't work.

No comments:

Post a Comment